# Evaluating diversion and treatment policies for opioid use disorder

### v3

@@ 3.3 Study Limitations
The model makes various assumptions based on data availability and to manage model complexity. The model does not generate individual characteristics except for age, which is used to estimate the time until a non-opioid-related death. The model incorporates some individual historical dependencies. The time in the inactive state depends on whether the individual came from the OUD treatment, CJS, hospital or ED, or active use state. Therefore, four separate distributions estimate the time in the inactive opioid use state. However, when individuals start using opioids (i.e., move to the active use state), they use the same distributions as stated in Table 1 to generate the next time until they are arrested, start OUD treatment, experience a hospital encounter, stop using opioids on their own, and experience an opioid-related death. Due to the sparsity of data, no other distributions take individual characteristics into account. As more data becomes available, future model versions could incorporate more individualistic data on risk and protective factors. To help obtain that level of data, this study could be used to justify additional data collection and information sharing.
@@
Due to a lack of criminal justice and treatment data differentiating between misused prescription opioids and illicitly manufactured opioids at the county level, we model opioid use generally. Therefore, opioid use and OUD are defined to encompass the effects of all opioids (e.g., prescription opioids, heroin, fentanyl) for all estimates. Similarly, we do not model seasonality; instead, we utilize and report yearly and cumulative estimates. Similarly, this study does not examine different types of OUD treatment and models OUD treatment in general. As discussed, various treatment options have various success rates and widespread use. As better treatments become more widely available, we expect even better outcomes as individuals stay inactive for longer.
@@
Individuals are assumed to complete their proven intervention successfully. However, in practice, individuals can be referred to a program but not complete it. In the context of an arrest, this is problematic since this would delay their time to the CJS state, staying in the active opioid use state for longer and leaving them exposed to additional risk of an opioid-related overdose, arrest, or death that is not accounted for in our current model. Additionally, individuals may continue to use opioids while in OUD treatment, which could increase opioid-related deaths and active opioid use from this study. However, those in treatment, especially if they are using MAT, have better outcomes that lower the risk of overdose and death than those not in treatment (Ma et al.,, 2019).
Lastly, we do not include the cost of implementing various diversions; we only include their cumulative savings of societal costs after policy implementation. We would expect this to reduce the overall savings both by a fixed amount and yearly operating costs that could be fairly substantial. For example, we do not account for the costs of OUD program coordinators in trying to connect individuals unsuccessfully to treatment programs. Instead, we model the societal costs, ignoring policy intervention implementation costs.
@@ 4 Conclusion
The results demonstrate that policies that create new pathways and programming by utilizing treatment services can lead to more opioid-resilient communities (e.g., communities with fewer adverse opioid-related outcomes). All three treatment policies can successfully create observable reductions in opioid-related arrests, hospital encounters, and opioid use when successfully diverting at least 20% of eligible individuals. These findings offer critical insights into the economic implications of different intervention strategies and can guide resource allocation and implementation decisions for jurisdictions. The three policies are shown to be complementary, with the largest impacts from overdose diversion, followed by case management and then arrest diversion. When all three are implemented, it could lead to potential 10-year cumulative societal savings of over \$1 Billion, less implementation costs.
@@ Acknowledgments
The authors would like to thank Captain Joseph Balles (retired) from the Madison Police Department and Dr. Aleksandra Zgierska, Professor of Family and Community Medicine, Public Health Sciences, and Anesthesiology and Perioperative Medicine at Penn State College of Medicine, for their subject matter expertise in providing information and feedback related to the data, and its analysis and modeling.
@@
The data and code that support the findings of this study are available at https://github.com/vmwhite/DES-Paper-Data-and-Analysis-2023. These data were derived from resources available in the public domain listed and cited in Table 1.
@@ A.1 Estimating Dane County Opioid Use Prevalence
Finally, equation (3) shows that for a given year, total opioid use prevalence can be estimated by the sum of Dane County heroin ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}^{\textit{DC,H}}$, and prescription opioid use, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}^{\textit{DC,P}}$, prevalence.
@@ A.4 Estimating Initiation and Prevalence Age
We estimated two separate distributions to sample individuals’ ages. Individuals generated in the starting population, described in Appendix A.2, use the prevalence age distribution. In contrast, individuals not in the starting population (i.e., during the warm-up and simulation periods) use the first initiation age distribution. Both age distributions use a log-normal distribution with a location parameter equal to 12 since we assume individuals are at least 12 years of age. We note that we first checked the NSDUH for opioid age of first use and the average age of individuals using opioids. However, the NSDUH for all years reports “nr,” i.e., not reported (SAMHSA, SAMHSA).
We estimate $\mu$ and $\sigma$ using the Australian study by Lewer et al., (2020) and equations (4)-(6). In Lewer et al., (2020), the figure entitled “2 Histograms of age at cohort entry” on page 9 of the appendix, we can estimate for the prevalence age distribution $m=30$ and $z_{0.75}=42.4$, resulting in $\mu=3.10$ and $\sigma=0.46$. The mode was selected by looking at the largest bar in the histogram. The first age distribution estimates the age when individuals enter the simulation through Distribution (1), i.e., the age at the time of initiation of opioid use. From the same figure in Lewer et al., (2020), we can estimate for the first initiation age distribution $m=25$ and $z_{0.75}=37.9$, resulting in $\mu=2.88$ and $\sigma=0.56$.
@@ A.5.2 Distribution (2) - Time in the active state given next event is an opioid-related death
Distribution (2) parameters are calculated separately for pre- and post-2019 due to the rapid increase in opioid-related deaths due to fentanyl in 2019 (Pardo et al.,, 2019). Distribution (2) is estimated via log-normal distributions with parameters $\mu_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{pre19}}}}$ and $\sigma_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{pre19}}}}$ for years before 2019 and $\mu_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{post19}}}}$ and $\sigma_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{post19}}}}$ for years 2019 and later. To estimate $\mu_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{pre19}}}}$ and $\mu_{2}^{\text{post19}}$, we use equations (4) - (6), where $\gamma=0$ and $m=1$ since individuals could experience a fatal overdose within the first day of use. We again assume $z_{q}=365.25$ days in a year. To estimate the respective quantiles, $q^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{pre19}}}}$ and $q^{\text{post19}}$ as the percentile of individuals that suffer an opioid-related death within one year, we first use the CDC wonder database and query for Dane County (National Center for Health Statistics,, 2022) to provide the number of deaths per year, i.e., $d_{i}^{\textit{DC}}$, and divide it by the prevalence of active opioid use in Dane County, i.e., $p^{DC}_{i}$ for all years $i$ that are available. Both the prevalence and death estimates are further described in Appendix A.4. Therefore, the quantiles represent the percentage of individuals with active opioid use who suffered an opioid-related death within the given year $i$ and can be calculated as:
@@
This results in $q^{\text{pre19}}=0.0027$ and $q^{\text{post19}}=0.0033$. We then estimated the log-normal distribution parameters using the equations (4) - (6), resulting in $\mu_{2}^{\text{pre19}}=17.60$, $\sigma_{2}^{\text{pre19}}=4.19$, $\mu_{2}^{\text{post19}}=17.13$, and $\sigma_{2}^{\text{post19}}=4.14$.
The parameters $\mu_{2}^{\text{pre19}}$ and $\mu_{2}^{\text{post19}}$ can be interpreted as the average number of log-normal days until the next opioid-related death per individual with active opioid use. Further, if an individual samples a time that is longer than the remaining length of the simulation, this can be interpreted that the individual will not suffer an opioid-related death. We note that ideally, we would directly estimate how long it takes for an individual to have an opioid-related death since most recently starting to use an opioid. However, this would be difficult to estimate as the vast majority of individuals who use opioids do not experience an opioid-related death (SAMHSA, SAMHSA). Therefore, estimating the time until opioid death only among those who experienced an opioid-related death would heavily skew the actual estimate among all individuals to the left. An alternative could be possible, such as only estimating an opioid-related death time for some percent of individuals, however, this would add additional parameters and variance to the model. In contrast, our estimation approach uses only four parameters, i.e., $\mu_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{pre19}}}}$, $\sigma_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{pre19}}}}$, $\mu_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{post19}}}}$, and $\sigma_{2}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{post19}}}}$ for a given set of years, which can estimate a time for all individuals.
@@ A.5.3 Distribution (3) - Time in the active state given next event is a hospital encounter
Distribution (3) is modeled as a log-normal distribution with parameters $\mu_{3}$ and $\sigma_{3}$ as specified in Table 1. To estimate $\mu_{3}$, we use the Wisconsin Department of Health Services, 2017b Hospital Encounter Data via the WISH Query. We query the rate for Dane County and ages 14+ since ages 12+ are not available. This data provides the number of hospital encounters $x_{i,3}$ for years $i=$ 2018 and 2019. We note that data from 2016 and 2017 were excluded during the calibration process to ensure the hospital encounter calibration targets were met. We then convert the data to reflect the mean number of days until the next hospital encounter per individual by multiplying the reciprocal by the number of days in the given year. We note that this estimates the days until the next hospital encounter among the active opioid use population rather than for an individual actively using opioids. We addressed this by dividing the mean number of days among the active use population by the prevalence of active opioid use in Dane County, or $p^{DC}_{i}$ for all years $i$ that are available. This estimate can be interpreted as the average number of days until the next opioid-related death per individual with active opioid use. Lastly, we average each yearly estimate and take the natural logarithm to obtain the final estimate for $\mu_{i,3}$. The calculation to estimate $\mu_{i,3}$ is:
@@ A.5.5 Distribution (5) - Time in the active state given next event is start OUD treatment
The distribution (5) parameters $\mu_{5}$ and $\sigma_{5}$ are estimated in the same way as the distribution (2) parameters. The Wisconsin DHS data provided the yearly number of OUD treatments, $x_{i,5}$ for $i=2016$ and $2017$. This resulted in $z_{0.063}=365.25$. We also use $\gamma=0$, and $m=610$ days. The most likely value $m$ is selected through model calibration by adding approximately $\pm$5from a starting value of 365. This resulted in $\mu_{5}=7.48$ and $\sigma_{5}=1.03$.

### v4

@@ 3.8 Study Limitations
The model makes various assumptions based on data availability and to manage model complexity. The model does not generate individual characteristics except for age, which is used to estimate the time until a non-opioid-related death. The model incorporates some individual historical dependencies. The time in the inactive state depends on whether the individual came from the OUD treatment, CJS, hospital or ED, or active use state. Therefore, four separate distributions estimate the time in the inactive opioid use state. However, when individuals start using opioids (i.e., move to the active use state), they use the same distributions as stated in Table 1 to generate the next time until they are arrested, start OUD treatment, experience a hospital encounter, stop using opioids on their own, and experience an opioid-related death. While the model accounts for the individual’s previous state, it assumes “memoryless” cumulative state durations that do not account for the cumulative effects of treatment, hospitalizations, and arrests. In reality, some state transitions may be history-dependent. For example, Nosyk et al., (2014) conclude that individuals who use heroin show cumulative time incarcerated delays cessation from using heroin, and Nosyk et al., (2009) conclude that subsequent treatment episodes of MAT with methadone tend to be longer in duration. Therefore, our model may overestimate active use to treatment enrollment for individuals with larger criminal histories while underestimating the duration of treatment for individuals in recurrent OUD treatment. We acknowledge there may be other cumulative impacts that are still being studied for individuals with opioid use that are not incorporated in this model. Due to the sparsity of data, no other distributions take individual characteristics into account. As more data becomes available, future model versions could incorporate more individualistic data on risk and protective factors. To help obtain that level of data, this study could be used to justify additional data collection and information sharing.
For modeling opioid-related deaths, we estimate county-level opioid use prevalence by redistributing the NSDUH state-level opioid use prevalence based on county-level overdose deaths. We acknowledge that the proportion of opioid use prevalence and opioid-related deaths likely do not exactly occur at consistent rates in the state of Wisconsin, which could impact the number of opioid-related deaths in Dane County estimated by the model. Additionally, NSDUH has been shown to underestimate the prevalence of drug use (Reuter et al.,, 2021). Therefore, the NSDUH data is another source of underestimation in the model when estimating both the amount of opioid use prevalence and opioid-related deaths since it relies on prevalence. This limitation underscores the importance of jurisdictions collecting accurate data regarding opioid use prevalence.
@@
Due to a lack of criminal justice and treatment data differentiating between misused prescription opioids and illicitly manufactured opioids at the county level, we model opioid use generally. Therefore, opioid use and OUD are defined to encompass the effects of all opioids (e.g., prescription opioids, heroin, fentanyl) for all estimates. Similarly, we do not model seasonality; instead, we utilize and report yearly and cumulative estimates. Additionally, this study does not examine different types of OUD treatment and models OUD treatment in general. As discussed, various treatment options have various success rates and widespread use. As better treatments become more widely available, we expect even better outcomes as individuals stay inactive for longer.
@@
Individuals are assumed to complete their proven intervention successfully. However, in practice, individuals can be referred to a program but not complete it. In the context of an arrest, this is problematic since this would delay their time to the CJS state, staying in the active opioid use state for longer and leaving them exposed to additional risk of an opioid-related overdose, arrest, or death that is not accounted for in our current model. Additionally, individuals may continue to use opioids while in OUD treatment. While those in treatment, especially those using MAT, have better outcomes that lower the risk of overdose and death than those not in treatment (Ma et al.,, 2019), a meta-analysis showed overdose mortality was 2.6 and 1.4 per 1000 persons in treatment using methadone and buprenorphine, respectively (Sordo et al.,, 2017). Therefore, our model may underestimate opioid-related deaths and active opioid use among individuals in OUD treatment.
Lastly, we do not include the cost of implementing various diversions; we only include their cumulative savings of societal costs after policy implementation. Additional fixed and variable cumulative policy implementation costs could be subtracted from the expected societal cost savings to get a total savings estimate that accounts for implementation.
@@ 4 Conclusion
The results demonstrate that policies that create new pathways and programming by utilizing treatment services can lead to more opioid-resilient communities (e.g., communities with fewer adverse opioid-related outcomes). All three treatment policies can successfully create observable reductions in opioid-related arrests, hospital encounters, and opioid use when successfully diverting at least 20% of eligible individuals. These findings offer critical insights into the economic implications of different intervention strategies and can guide resource allocation and implementation decisions for jurisdictions. The three policies are shown to be complementary, with the largest impacts from overdose diversion, followed by case management, and then arrest diversion, assuming the same level of implementation increase in all policies. When all three are implemented, it could lead to potential 10-year cumulative societal savings of over \$500 million, less implementation costs, which we estimated to be approximately one tenth of the savings of the case study’s arrest diversion program.
@@ Acknowledgments
The authors would like to thank Captain Joseph Balles (retired) from the Madison Police Department and Dr. Aleksandra Zgierska, Professor of Family and Community Medicine, Public Health Sciences, and Anesthesiology and Perioperative Medicine at Penn State College of Medicine, for their subject matter expertise in providing information and feedback related to the data, and its analysis and modeling. The authors would like to thank the anonymous reviewers and the associate editor for their valuable comments and suggestions to improve the quality of the paper.
@@
The data and code that support the findings of this study are available at https://doi.org/10.5281/zenodo.15498919. These data were derived from resources available in the public domain listed and cited in Table 1.
@@ A.1 Estimating Dane County Opioid Use Prevalence
Finally, equation (3) shows that for a given year, total opioid use prevalence can be estimated by the sum of Dane County heroin ${p}^{\textit{DC,H}}$, and prescription opioid use, ${p}^{\textit{DC,P}}$, prevalence.
@@ A.4 Estimating Initiation and Prevalence Age
We estimated two separate distributions to sample individuals’ ages. Individuals generated in the starting population, described in Appendix A.2, use the prevalence age distribution. In contrast, individuals not in the starting population (i.e., during the warm-up and simulation periods) use the first initiation age distribution. Both age distributions use a log-normal distribution with a location parameter equal to 12 since we assume individuals are at least 12 years of age. We also assume a maximum enter age of 105 years; this is done by truncating the age log-normal distributions at 105 years by resampling until a value below 105 years is generated. We estimate $\mu$ and $\sigma$ using SAMHSA (2019) and equations (4)-(6). From SAMHSA (2019), we use Table 1.19A to estimate the prevalence age distribution. We can estimate mode $m=32$ and quantile ${x}_{0.933}=65$, resulting in $\mu=3.74$ and $\sigma=0.49$. The mode was selected by taking the midpoint of the age range with the largest population percentage (e.g., $\frac{30+34}{2}=32$) in 2018. The quartile was selected by summarizing the percentage of all age groups below 65 years old, of which “65 and older” is the maximum age reported in Table 1.19A. The year 2018 was selected as it had the largest percentile of underlying populations between reported years. Using similar logic, we use Tables 4.5, 4.7, and 4.8 to estimate the initial age of individuals when they enter the simulation through Distribution (1), i.e., the age at the time of initiation of opioid use. We can estimate for the first initiation age distribution using $m=21.5$ and ${x}_{0.403}=26$ in 2019, resulting in $\mu=2.08$ and $\sigma=0.76$.
@@ A.5.2 Distribution (2) - Time in the active state given next event is an opioid-related death
Distribution (2) parameters are calculated separately for pre- and post-2019 due to the rapid increase in opioid-related deaths due to fentanyl in 2019 (Pardo et al.,, 2019). Distribution (2) is estimated via log-normal distributions with parameters $\mu_{2}^{{{\text{pre19}}}}$ and $\sigma_{2}^{{{\text{pre19}}}}$ for years before 2019 and $\mu_{2}^{{{\text{post19}}}}$ and $\sigma_{2}^{{{\text{post19}}}}$ for years 2019 and later. To estimate $\mu_{2}^{{{\text{pre19}}}}$ and $\mu_{2}^{\text{post19}}$, we use equations (4) - (6), where $\gamma=0$ and $m=1$ since individuals can experience a fatal overdose within the first day of use. We again assume ${x}_{q}=365.25$ days in a year. To estimate the respective quantiles, $q^{{{\text{pre19}}}}$ and $q^{\text{post19}}$ as the percentile of individuals that suffer an opioid-related death within one year, we first use the CDC wonder database and query for Dane County (National Center for Health Statistics,, 2022) to provide the number of deaths per year, i.e., $d_{i}^{\textit{DC}}$, and divide it by the prevalence of active opioid use in Dane County, i.e., $p^{DC}_{i}$ for all years $i$ that are available. Both the prevalence and death estimates are further described in Appendix A.1. Therefore, the quantiles represent the percentage of individuals with active opioid use who suffered an opioid-related death within a year and can be calculated as:
@@
This results in $q^{\text{pre19}}=0.0027$ and $q^{\text{post19}}=0.0033$. We then estimated the log-normal distribution parameters using the equations (4) - (6), resulting in $\mu_{2}^{\text{pre19}}=17.60$, $\sigma_{2}^{\text{pre19}}=4.19$, $\mu_{2}^{\text{post19}}=17.13$, and $\sigma_{2}^{\text{post19}}=4.14$. We note that a very small value of $q$ is estimated in (7); which may limit our study by underestimating the variability in individuals’ time to opioid-related death. As a result, the simulated times may be more clustered or have a narrower range than what we would observe in the real world. However, since our evaluation focuses on population-level outcomes rather than individual trajectories, we believe the underestimation would have a limited impact on our results and conclusions.
To aid the interpretation of the quantile, consider the annual overdose death rate among the active use population, $\nicefrac{{d^{DC}_{2019}}}{{p^{DC}_{2019}}}$. The reciprocal of $\nicefrac{{d^{DC}_{2019}}}{{p^{DC}_{2019}}}$ can be interpreted as the average time until overdose death, since the start of 2019. When we average this over multiple years, the reference point of “since the start of the year” becomes an arbitrary year. We instead interpret “starting active opioid use” as the beginning of the time period. This yields an estimate of time-to-overdose death from the onset of use. However, this estimate represents a lower bound, as the prevalence data likely underestimates the true number of active users.
The parameters $\mu_{2}^{\text{pre19}}$ and $\mu_{2}^{\text{post19}}$ can be interpreted as the average number of log-normal days until the next opioid-related death per individual with active opioid use. Further, if an individual samples a time that is longer than the remaining length of the simulation, this can be interpreted that the individual will not suffer an opioid-related death. We note that ideally, we would directly estimate how long it takes for an individual to have an opioid-related death since most recently starting to use an opioid. However, this would be difficult to estimate as the vast majority of individuals who use opioids do not experience an opioid-related death (SAMHSA, SAMHSA). Therefore, estimating the time until opioid death only among those who experienced an opioid-related death would heavily skew the actual estimate among all individuals to the left. An alternative could be possible, such as only estimating an opioid-related death time for some percent of individuals, however, this would add additional parameters and variance to the model. In contrast, our estimation approach uses only four parameters, i.e., $\mu_{2}^{{{\text{pre19}}}}$, $\sigma_{2}^{{{\text{pre19}}}}$, $\mu_{2}^{{{\text{post19}}}}$, and $\sigma_{2}^{{{\text{post19}}}}$ for a given set of years, which can estimate a time for all individuals.
@@ A.5.3 Distribution (3) - Time in the active state given next event is a hospital encounter
Distribution (3) is modeled as a log-normal distribution with parameters $\mu_{3}$ and $\sigma_{3}$ as specified in Table 1. To estimate $\mu_{3}$, we use the Wisconsin Department of Health Services, 2017b Hospital Encounter Data via the WISH Query. We query the rate for Dane County and ages 14+ since ages 12+ are not available. This data provides the number of hospital encounters $x_{i,3}$ for years $i=$ 2018 and 2019. We note that data from 2016 and 2017 were excluded during the calibration process to ensure the hospital encounter calibration targets were met. We then convert the data to reflect the mean number of days until the next hospital encounter among the active opioid use population by multiplying the reciprocal by the number of days in the given year, $D_{i}$. To obtain a $\mu_{3}$ that estimates the days until the next hospital encounter for the average individual actively using opioids, rather than among the active opioid use population, we multiply by the prevalence of active opioid use in Dane County, or $p^{DC}_{i}$ for all years $i$ that are available. Lastly, we average each yearly estimate and take the natural logarithm to obtain the final estimate for $\mu_{i,3}$. The calculation to estimate $\mu_{i,3}$ is:
@@
For example, if the average number of days until the next hospital encounter is 40 days (i.e., $D_{i}/x_{i,3}=40$), then the mean number of days for an individual’s next hospital encounter is 400 days (i.e., $D_{i}p_{i}/x_{i,3}=4000$), or more than 10 years. As discussed in Section A.5.2, most individuals who use opioids will not experience an opioid-related death within the simulation run time. The evidence is similar for hospital encounters, opioid-related arrests, and OUD treatment starts. Therefore, if an individual is scheduled for a hospital encounter time beyond the end of the simulation, this would be interpreted as they did not experience a hospital encounter. Justification for this modeling choice can be seen in Table 3 by comparing the estimated prevalence of 2017 opioid use against 2017 opioid-related deaths, hospital encounters, opioid-related arrests, and OUD treatment starts.
@@ A.5.5 Distribution (5) - Time in the active state given next event is start OUD treatment
The distribution (5) parameters $\mu_{5}$ and $\sigma_{5}$ are estimated in the same way as the distribution (2) parameters. The Wisconsin DHS data provided the yearly number of OUD treatments, $x_{i,5}$ for $i=2016$ and $2017$. This resulted in ${x}_{0.063}=365.25$. We also use $\gamma=0$, and $m=610$ days. The most likely value $m$ is selected through model calibration by adding approximately $\pm$5from a starting value of 365. This resulted in $\mu_{5}=7.48$ and $\sigma_{5}=1.03$.